OpenStudy (anonymous):

find the limit as x approaches 0 for 3x(cos 3x-1)/((sin 2x)-2x)

OpenStudy (anonymous):

\[\lim_{x \rightarrow o} \frac{ 3x \cos (3x-1) }{ \sin 2x -2x}\]

OpenStudy (anonymous):

no

OpenStudy (anonymous):

because I re-wrote the equation so I could see if anyone knew how to solve it

OpenStudy (anonymous):

yeah.....

OpenStudy (experimentx):

use L'hopital rule ...

OpenStudy (anonymous):

how many times?

OpenStudy (anonymous):

I have tried that multiple times

OpenStudy (experimentx):

woops!! looks like i misunderstood question.

OpenStudy (experimentx):

http://www.wolframalpha.com/input/?i=limit+x-%3E0+3x%28cos+3x-1%29%2F%28%28sin+2x%29-2x%29 definitely L'hopital rule will give you the required limit.

OpenStudy (experimentx):

 use Taylor expansion after ... do you know how to use it?